# Cost-Efficient Online Decision Making: A Combinatorial Multi-Armed Bandit Approach

### v2

@@ 1 Introduction
We take a novel view of the online decision making problem described above, where we cast it as a combinatorial multi-armed bandit problem (CMAB) (Cesa-Bianchi and Lugosi, 2012), an extension of the classical multi-armed bandit problem (MAB). The standard MAB problem is a common way of representing the trade-off inherent to decision making problems between exploring an environment by performing successive actions (playing arms, in bandit nomenclature) to gain more knowledge and exploiting that knowledge to reach a long-term objective. In a CMAB problem, an agent interacts with the environment in each time step by performing a set of actions (a super arm), where the set may be subject to combinatorial constraints. Each individual action (base arm) in that set may result in some observed feedback (so-called semi-bandit feedback), and the agent also receives a reward associated with the complete super arm.
[…]
@@
    We study the problem of cost-efficient online decision making, where costs may be stochastic and depend on both test and decision outcomes. For this, we propose a novel combinatorial semi-bandit framework, providing a new perspective on cost-efficient information acquisition.
@@
[…]
@@ 2 Related Work
The primary goal of cost-efficient information acquisition algorithms is to sequentially select from a number of available costly tests to make a decision (such as making a prediction about a label) with a minimal cost. To minimize cost, these tests are performed until a sufficient level of confidence in the decision is reached. Prominent algorithms for cost-efficient information acquisition are Information Gain (IG) (Zheng et al., 2012), EC2 (Golovin et al., 2010), and Uncertainty Sampling (US). Among them, EC2 has been proved to enjoy adaptive submodularity (Golovin et al., 2010), and thus yields near-optimal cost. The works of Rahbar et al. (2023); Chen et al. (2017) consider active information acquisition in an online setting. The regret bounds in these works are based on formulating the problem as a Partially Observable Markov Decision Process (POMDP) and are exponential in the total number of possible tests. In this work, we consider a CMAB formulation of the problem and derive a regret bound that is linear in the number of tests (and sublinear w.r.t. the horizon $T$). In addition, unlike Rahbar et al. (2023); Chen et al. (2017), our framework supports variability in the cost of the tests depending on the outcomes of the tests and the decisions.
@@
CMAB methods have been utilized to address online learning problems in various settings but often without considering cost-efficient information acquisition. Durand and Gagné (2014) adapt a combinatorial variant of Thompson Sampling for online feature selection where the agent has a fixed budget for tests and the reward is a linear function of the test outcomes. Both assumptions are used in later works studying the problem from a CMAB perspective (Bouneffouf et al., 2017, 2021), but differ from the setting considered in this work (e.g., the budget for information acquisition is fixed with no concept of achieving a decision, and the costs are independent of the outcomes of the tests and the decisions). Åkerblom et al. (2023) analyze the Bayesian regret of combinatorial Thompson Sampling using a similar technique as this work (originally developed by Russo and Van Roy (2014) for standard and linear MAB), but for a very different super arm reward function and setting (for the online minimax path problem, which does not consider cost-efficient information acquisition and decision making).
@@ 3 Problem Formulation
In this section, we propose a new formulation of the online decision making problem which incorporates the cost of acquiring test outcomes. In each time step $t$, we receive the problem instance (e.g., data point) $\mathbf{x}^{(t)}$ with $n$ tests $\{x_{1}^{(t)},x_{2}^{(t)},\dots,x_{n}^{(t)}\}$. Here, we assume that all test outcomes (unknown until after each test is performed) are binary (i.e., $x_{i}^{(t)}\in\{0,1\}$ for all $i$), but our methods can easily be extended to other test values (see Section 6). The goal is to make an accurate decision $y^{(t)}\in[m]$ for that problem instance with a minimal cost of performing tests ($m$ represents the total number of options for the decision, or outcomes if the correct decision is viewed as a random variable), where we denote the correct decision by $y^{*}$.
@@
Consistent with the common Naïve Bayes assumption, we let all tests $x_{i}^{(t)}$ be mutually independent and Bernoulli-distributed conditional on the decision $y^{(t)}$. We also let $y^{*}$ be known given a full realization of $\mathbf{x}^{(t)}$. This implies that upon performing all available tests, we know the correct decision for the problem instance. This is realistic and consistent with the settings of Golovin et al. (2010); Chen et al. (2017); Rahbar et al. (2023), though they do not follow a CMAB formulation. As an example, consider a medical diagnosis problem. This property means that if we perform all available medical tests, we can determine the correct diagnosis.
Regarding the costs, we assume that they directly depend on the outcomes of the tests and the decisions. As motivation, consider, for example, a medical test to confirm the presence of some indicator for a particular disease. If such an indicator is found, the test can be terminated immediately with a positive outcome, whereas establishing the non-presence of an indicator (for a negative test outcome) may require more time and be more invasive.
@@
We model the problem described above by an elegant design of a stochastic combinatorial multi-armed bandit (CMAB) problem (Cesa-Bianchi and Lugosi, 2012) with semi-bandit feedback. Briefly, CMAB corresponds to a reward maximization problem with a set of base arms. In each time step, the agent selects a subset of base arms. After playing this subset of base arms, called a super arm, the agent receives a reward from the environment. The objective is to maximize the expected sum of rewards within a time horizon $T$, which is typically formulated as a regret minimization task.
@@
In our problem, since the costs of tests depend on both test and decision outcomes, we need to consider a set of base arms that reflect both. Therefore, we define the set $\mathcal{A}$ of $n\times m$ base arms $a_{ij}$ ($i\in[n]$ and $j\in[m]$). In each time step $t$, the agent selects a super arm $S^{(t)}\in 2^{\mathcal{A}}$ and receives the reward $R(S^{(t)})=R^{(t)}$. A base arm $a_{ij}$ in $S^{(t)}$ indicates that the test with index $i$ is performed for making decision $y^{(t)}=j$. We define $\mathcal{I}^{(t)}\subseteq 2^{\mathcal{A}}$ as the set of feasible super arms at time $t$ which yield a decision, i.e., each feasible super arm consists of tests and a single decision (concretely, if $a_{ij},a_{lk}\in S\in\mathcal{I}^{(t)}$, then $j=k$). The feedback at time $t$ of a base arm $a_{ij}\in\mathcal{A}$ (observable by the agent if $a_{ij}\in S^{(t)}$) is the corresponding outcome of test $x_{i}^{(t)}$, which given a decision $y^{(t)}=j$ is Bernoulli-distributed with an unknown (to the agent) parameter $\theta_{ij}\triangleq Pr(x_{i}=1|y=j)$, for all $i,j$. While tests may, in practice, be selected sequentially, the super arm (and base arms) are only known in hindsight after the decision has been made. This is intentional and allows us to analyze the problem as a CMAB.
@@
We let the cost (assumed to be fixed and known) of each base arm $a_{ij}$ be $\mu_{ij}^{(0)}\in[0,1]$ and $\mu_{ij}^{(1)}\in[0,1]$ (given test outcomes 00 and $1$, respectively). Then, we define the expected reward of the selected feasible super arm $S^{(t)}$ as the negative sum of expected base arm costs, such that $\mathbb{E}[R(S^{(t)})]=-\sum_{ij\in S}\mu_{ij}$, where
@@
with the vector of all $\mu_{ij}$’s denoted by $\boldsymbol{\mu}$ and the vector of all $\theta_{ij}$’s denoted by $\boldsymbol{\theta}$. We let $\boldsymbol{\mu}^{*}$ be the true underlying mean cost vector, with corresponding parameter vector $\boldsymbol{\theta}^{*}$. We define the suboptimality gap at time $t$ of a feasible super arm $S$ as
@@ 4 CMAB Methods for ODM
The $\theta_{ij}^{*}$ values of the true parameter vector $\boldsymbol{\theta}^{*}$ are initially unknown, but to be able to employ prior knowledge, we consider a prior distribution over $\theta_{ij}^{*}$’s. Formally, we assume that each $\theta_{ij}^{*}$ follows a Beta distribution, $\mathrm{Beta}(\alpha_{ij}^{(0)},\beta_{ij}^{(0)})$. We also assume a fixed and known distribution over decision outcomes, denoted by $P(y)$ (i.e., the learning agent knows the marginal distribution a priori). However, even if the oracle is assumed to always be able to find the correct decision (i.e., it only needs to minimize the number of tests), the optimization problem is NP-hard, and thus in practice, we use approximate oracles (introduced in Section 4.1).
@@
Algorithm 1 summarizes our framework, with Thompson Sampling used as exploration method. In each time step $t$, we sample the parameters from the current posterior distribution (line 2). We then (in lines 3-4) use the oracle to get a super arm based on the sampled parameters, and perform the tests sequentially. Finally, we use Algorithm 2 to update the posterior distribution of the parameters. Algorithm 2 uses the fact that Beta distribution is conjugate prior to the Bernoulli distribution [^1] Since test outcomes are binary, then, given decision $y=j$, each test $x_{i}$ follows a Bernoulli distribution with parameter $\theta_{ij}$.
@@ 4.1 Cost-Efficient Approximate Oracles
Before introducing W-IG and W-EC2, we need to define the notions of hypothesis and decision region. We call a full realization of test outcomes (i.e., of all possible tests) a hypothesis. Therefore, with $n$ binary tests, we have $2^{n}$ different possible hypotheses. Let $\mathcal{H}\triangleq\{0,1\}^{n}$ be the set of all possible hypotheses for $n$ tests. We can partition $\mathcal{H}$ into $m$ disjoint sets. We call each of these disjoint sets a decision region. Each decision region corresponds to a decision $y\in[m]$. So our objective will be finding the correct decision region, for which the cost of performing tests is low. We want to emphasize that one important difference between these oracles and standard CMAB oracles is that they are interactive. In other words, during a single time step, they sequentially perform tests and observe outcomes to select subsequent tests and make a decision.
@@
Also like with W-IG, in the W-EC2 algorithm, we sequentially perform tests that maximize $\Delta_{\text{W-EC}^{2}}\left(i\mid\mathbf{x}_{\mathcal{P}}\right)$. We stop performing tests when all edges are cut, meaning that there is only one decision region left.

### v3

# algorithmic

@@ 1 Introduction
We take a novel view of the online decision making problem described above, where we cast it as a combinatorial multi-armed bandit problem (CMAB) (Cesa-Bianchi & Lugosi, 2012), an extension of the classical multi-armed bandit problem (MAB). The standard MAB problem is a common way of representing the trade-off inherent to decision making problems between exploring an environment by performing successive actions (playing arms, in bandit nomenclature) to gain more knowledge and exploiting that knowledge to reach a long-term objective. In a CMAB problem, an agent interacts with the environment in each time step by performing a set of actions (a super arm), where the set may be subject to combinatorial constraints. Each individual action (base arm) in that set may result in some observed feedback (so-called semi-bandit feedback), and the agent also receives a reward associated with the complete super arm.
In this work, we view the combined choice of tests to perform and the selected decision as a super arm in a CMAB problem where the underlying cost distribution parameters are initially unknown, which allows us to leverage effective MAB exploration strategies like Thompson Sampling (TS) (Thompson, 1933) and Upper Confidence Bound (UCB) (Auer et al., 2002) together with cost-efficient (active) information acquisition methods, such as Equivalent Class Edge Cutting (EC2) (Golovin et al., 2010) and Information Gain (IG) (Dasgupta, 2005). We summarize our contributions as follows.
@@
    We study the novel problem of outcome-dependent cost-efficient online decision making, where costs may be stochastic and depend on both test and decision outcomes.
@@
    We propose a novel combinatorial semi-bandit framework for our online decision making problem. This formulation involves an elegant definition of base arms and super arms, providing a new perspective on cost-efficient information acquisition while simultaneously exploring the concept of interactive CMAB oracles.
@@
    We adapt a number of bandit methods to this framework, namely Thompson Sampling and Upper Confidence Bound, combined with novel extensions of EC2 and IG which can handle tests with different and stochastic costs, which we call W-EC2 and W-IG, respectively. Thereby, in addition to Thompson Sampling, we develop W-EC2 and W-IG oracles utilizing a Bayesian variant of UCB (called BayesUCB by Kaufmann et al. 2012) that fits well to our framework and enables the UCB method to effectively employ prior knowledge in the form of a prior distribution.
  4. 4.
@@ 2 Related Work
The primary goal of cost-efficient information acquisition (also called Bayesian active learning in the literature) algorithms is to sequentially select from a number of available costly tests to make a decision (such as making a prediction about a label) with a minimal cost. To minimize cost, these tests are performed until a sufficient level of confidence in the decision is reached. Prominent algorithms for cost-efficient information acquisition are Information Gain (IG) (originally developed by Lindley 1956), EC2 (Golovin et al., 2010), and Uncertainty Sampling (US). Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011) proposes an equivalent formulation of the underlying utility function in IG which provides practical and computational advantages. EC2 has been proved to enjoy adaptive submodularity (Golovin et al., 2010), and thus yields near-optimal cost. Acquisition functions in Bayesian optimization have also been used in active learning. Probability of Improvement (PI) (Kushner, 1964) and Expected Improvement (EI) (Jones et al., 1998) are two important acquisition functions used in active learning. Both PI and EI use surrogate models (e.g., Gaussian Processes) to select the next sample to query (see Di Fiore et al. (2024) for more details). The works of Rahbar et al. (2023); Chen et al. (2017) consider active information acquisition in an online setting. The regret bounds in these works are based on formulating the problem as a Partially Observable Markov Decision Process (POMDP) and are exponential in the total number of possible tests. In this work, we consider a CMAB formulation of the problem and derive a regret bound that is linear in the number of tests (and sublinear w.r.t. the horizon $T$). In addition, unlike Rahbar et al. (2023); Chen et al. (2017), our framework supports variability in the cost of the tests depending on the outcomes of the tests and the decisions.
@@
CMAB methods have been utilized to address online learning problems in various settings but often without considering cost-efficient information acquisition. Durand & Gagné (2014) adapt a combinatorial variant of Thompson Sampling for online feature selection where the agent has a fixed budget for tests and the reward is a linear function of the test outcomes. Both assumptions are used in later works studying the problem from a CMAB perspective (Bouneffouf et al., 2017; 2021), but differ from the setting considered in this work (e.g., the budget for information acquisition is fixed with no concept of achieving a decision, and the costs are independent of the outcomes of the tests and the decisions). Wang & Chen (2018) analyze the regret of combinatorial Thompson sampling, but their proof technique is not directly applicable to our problem since they do not allow changes in the set of available super arms. Therefore, in this work, we adopt the approach originally developed by Russo & Van Roy (2014) for standard and linear MAB. Finally, it is notable that the analysis of Wang & Chen (2018) yields an instance dependent regret bound, whereas our analysis is focused on deriving Bayesian and instance-independent bounds.
Online decision making.
Our work relates to the theoretical online decision making framework presented by Cesa-Bianchi et al. (2021). Their framework aims to maximize total Return on Investment (ROI) in a sequence of decision making tasks, where each task involves accepting or rejecting an innovation. The authors provide an algorithm to learn ROI-maximizing decision making policies, with theoretical guarantees of convergence to an optimal policy. In the framework, accept/reject decisions are made by drawing i.i.d. samples from a probability distribution modeling the value of an innovation. In contrast, our approach involves making low-cost decisions by performing a sequence of different tests (with different costs). Furthermore, our work is not limited to binary accept/reject decisions.
Algorithm 1 Online decision making protocol
[…]
2:  for t = 1, 2, …, T do
3:     Receive problem instance $\mathbf{x}^{(t)}$.
4:     for all selected tests $x_{i}^{(t)}$ (by an oracle) do
5:        Observe outcome of test $x_{i}^{(t)}\sim\text{Bernoulli}(\theta_{ij})$.
6:     end for
7:     Make a decision $y^{(t)}=j\in[m]$ based on test outcomes.
8:     Super arm $S^{(t)}=\{a_{ij}\in\mathcal{A}|\text{test }x_{i}^{(t)}\text{is performed}\}\in\mathcal{I}^{(t)}$ where $\mathcal{I}^{(t)}\subseteq 2^{\mathcal{A}}$ is set of feasible super arms.
9:     Receive reward $R(S^{(t)})$ where $\mathbb{E}[R(S^{(t)})]=-\sum_{ij\in S^{(t)}}\mu_{ij}$ and $\mu_{ij}=\mu_{ij}^{(1)}\times\theta_{ij}+\mu_{ij}^{(0)}\times(1-\theta_{ij})$.
10:  end for
@@ 3 Problem Formulation
In this section, we propose our new formulation of the online decision making problem which incorporates the cost of acquiring test outcomes. In each time step $t$, we receive the problem instance (e.g., data point) $\mathbf{x}^{(t)}$ with $n$ tests $\{x_{1}^{(t)},x_{2}^{(t)},\dots,x_{n}^{(t)}\}$. Here, we assume that all test outcomes (unknown until after each test is performed) are binary (i.e., $x_{i}^{(t)}\in\{0,1\}$ for all $i$), but our methods can easily be extended to other test values (see Section 6.3). The goal is to make an accurate decision $y^{(t)}\in[m]$ for that problem instance with a minimal cost of performing tests ($m$ represents the total number of options for the decision, or outcomes if the correct decision is viewed as a random variable), where we denote the correct decision by $y^{*}$. As an example, let’s consider the medical problem of diagnosing lung cancer. At each time step $t$, a patient $\mathbf{x}^{(t)}$ arrives, and our goal is to make a decision (prediction) $y^{(t)}\in\{1(benign),2(malignant),3(infection)\}$ ($m=3$). In this problem, there are several tests that can be performed to reach a decision. For instance, available tests can include a CT [^1] Computed Tomography scan, a PET222 Positron Emission Tomography scan, genetic testing and a biopsy ($n=4$). These tests can have positive or negative outcomes (test results), which will be used in the decision making process.
@@
Consistent with the common Naïve Bayes assumption, we let all tests $x_{i}^{(t)}$ be mutually independent and Bernoulli-distributed conditional on the decision $y^{(t)}$. The Naïve Bayes assumption is common for sequential decision making, even in the offline setting, and is consistent with previous works including Chen et al. (2017); Rahbar et al. (2023). This assumption is needed for computing the gains in our approximate oracles defined in Section 4.1 (W-IG and W-EC2). We also let $y^{*}$ be known given a full realization of $\mathbf{x}^{(t)}$. This implies that upon performing all available tests, we know the correct decision for the problem instance. This is realistic and consistent with the settings of Golovin et al. (2010); Chen et al. (2017); Rahbar et al. (2023), though they do not follow a CMAB formulation. As an example, consider a medical diagnosis problem. This property means that if we perform all available medical tests, we can determine the correct diagnosis.
Regarding the costs, we assume that they directly depend on the outcomes of the tests and the decisions. As motivation, again consider the medical diagnosis problem. Here, the costs (e.g., time and patient’s discomfort) of medical tests can vary based on the underlying affliction or the outcome of the test itself. For example, in the case of diagnosing cancer through biopsy procedures, the cost of the biopsy can change based on the specific underlying condition and the outcome of the test. If the suspected cancerous tissue can be accessed easily and shows a positive result for cancer cells, the biopsy procedure may cause lower costs in terms of time and discomfort. Conversely, if the tissue is in a challenging area or requires multiple biopsies, the procedure’s complexity may cause higher costs, both in terms of time and increased patient’s discomfort.
@@
We model the problem described above by an elegant design of a stochastic combinatorial multi-armed bandit (CMAB) problem (Cesa-Bianchi & Lugosi, 2012) with semi-bandit feedback. Briefly, CMAB corresponds to a reward maximization problem with a set of base arms. In each time step, the agent selects a subset of base arms. After playing this subset of base arms, called a super arm, the agent receives a reward from the environment. The objective is to maximize the expected sum of rewards within a time horizon $T$, which is typically formulated as a regret minimization task.
Figure 1: Illustration of base arms and super arms in our problem formulation. In this example, we have four tests ($n=4$) and two possible decisions ($m=2$). Based on these tests and decisions, we have $4\times 2$ possible base arms $\mathcal{A}=\{a_{ij}|i\in[4],j\in[2]\}$. If we perform tests with indices $2,4,1$, and we make decision $y=2$ (based on the test results), then the super arm will be $S=\{a_{22},a_{42},a_{12}\}$.
@@
In our problem, since the costs of tests depend on both test and decision outcomes, we need to consider a set of base arms that reflect both. Therefore, we define the set $\mathcal{A}$ of $n\times m$ base arms $a_{ij}$ ($i\in[n]$ and $j\in[m]$). In each time step $t$, the agent selects a super arm $S^{(t)}\in 2^{\mathcal{A}}$ and receives the reward $R(S^{(t)})=R^{(t)}$. A base arm $a_{ij}$ in $S^{(t)}$ indicates that the test with index $i$ is performed for making decision $y^{(t)}=j$ (see Figure 1). We define $\mathcal{I}^{(t)}\subseteq 2^{\mathcal{A}}$ as the set of feasible super arms at time $t$ which yield a decision, i.e., each feasible super arm consists of tests and a single decision (concretely, if $a_{ij},a_{lk}\in S\in\mathcal{I}^{(t)}$, then $j=k$). The feedback at time $t$ of a base arm $a_{ij}\in\mathcal{A}$ (observable by the agent if $a_{ij}\in S^{(t)}$) is the corresponding outcome of test $x_{i}^{(t)}$, which given a decision $y^{(t)}=j$ is Bernoulli-distributed (test outcomes are binary) with an unknown (to the agent) parameter $\theta_{ij}\triangleq Pr(x_{i}=1|y=j)$, for all $i,j$. While tests may, in practice, be selected sequentially, the super arm (and base arms) are only known in hindsight after the decision has been made. This is intentional and allows us to analyze the problem as a CMAB.
@@
We let the cost (assumed to be fixed and known) of each base arm $a_{ij}$ be $\mu_{ij}^{(0)}\in[0,1]$ and $\mu_{ij}^{(1)}\in[0,1]$ (given test outcomes 00 and $1$, respectively). Then, we define the expected reward of the selected feasible super arm $S^{(t)}$ as the negative sum of expected base arm costs, such that $\mathbb{E}[R(S^{(t)})]=-\sum_{ij\in S^{(t)}}\mu_{ij}$, where
@@
with the vector of all $\mu_{ij}$’s denoted by $\boldsymbol{\mu}$ and the vector of all $\theta_{ij}$’s denoted by $\boldsymbol{\theta}$. We let $\boldsymbol{\mu}^{*}$ be the true underlying mean cost vector, with corresponding parameter vector $\boldsymbol{\theta}^{*}$. This is a novel problem; the previous studies are special cases of this problem with $\mu_{ij}^{(0)}=\mu_{ij}^{(1)}=c_{i}$, where $c_{i}$ is the cost of performing test $i$ and is independent of test outcome or the decision. We show an outline of our online decision making protocol in Algorithm 1.
We define the suboptimality gap at time $t$ of a feasible super arm $S$ as
@@ 4 CMAB Methods for ODM
The $\theta_{ij}^{*}$ values of the true parameter vector $\boldsymbol{\theta}^{*}$ are initially unknown, but to be able to employ prior knowledge, we consider a prior distribution over $\theta_{ij}^{*}$’s. Formally, we assume that each $\theta_{ij}^{*}$ follows a Beta distribution, $\mathrm{Beta}(\alpha_{ij}^{(0)},\beta_{ij}^{(0)})$. The reason is that a Beta distribution is defined on range $[0,1]$, and parameter of a Bernoulli distribution should be in this range (each $\theta_{ij}^{*}$ is parameter of a Bernoulli distribution). We also assume a fixed and known distribution over decision outcomes, denoted by $P(y)$ (i.e., the learning agent knows the marginal distribution a priori). However, even if the oracle is assumed to always be able to find the correct decision (i.e., it only needs to minimize the number of tests), the optimization problem is NP-hard, and thus in practice, we use approximate oracles (introduced in Section 4.1).
@@
Algorithm 2 summarizes our framework, with Thompson Sampling used as exploration method. In each time step $t$, we sample the parameters from the current posterior distribution (line 2). We then (in lines 3-4) use the oracle to get a super arm based on the sampled parameters, and perform the tests sequentially. Finally, we use Algorithm 3 to update the posterior distribution of the parameters. Algorithm 3 uses the fact that Beta distribution is conjugate prior to the Bernoulli distribution. [^3] Since test outcomes are binary, then, given decision $y=j$, each test $x_{i}$ follows a Bernoulli distribution with parameter $\theta_{ij}$.
@@ 4.1 Cost-Efficient Approximate Oracles
Before introducing W-IG and W-EC2, we need to define the notions of hypothesis and decision region. We call a full realization of test outcomes (i.e., of all possible tests) a hypothesis. Therefore, with $n$ binary tests, we have $2^{n}$ different possible hypotheses. Let $\mathcal{H}\triangleq\{0,1\}^{n}$ be the set of all possible hypotheses for $n$ tests. We can partition $\mathcal{H}$ into $m$ disjoint sets. We call each of these disjoint sets a decision region. Each decision region corresponds to a decision $y\in[m]$. So our objective will be finding the correct decision region, for which the cost of performing tests is low. We want to emphasize that one important difference between these oracles and standard CMAB oracles is that they are interactive. In other words, during a single time step, they sequentially perform tests and observe outcomes to select subsequent tests and make a decision. To our knowledge, our work is the first work which utilizes this approach for CMABs.
@@
Also like with W-IG, in the W-EC2 algorithm, we sequentially perform tests that maximize the gain $\Delta_{\text{W-EC}^{2}}\left(i\mid\mathbf{x}_{\mathcal{P}}\right)$. We stop performing tests when all edges are cut, which means that we have only one decision region left.